OpenStudy (ragingsquirrel):

The base of a rectangular prism is 20 cm2. If the volume of the prism is 100 cm3, what is its height? 5 cm 10 cm 80 cm 120 cm

OpenStudy (ragingsquirrel):

A full explanation would be good here.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

area*height=volume

OpenStudy (anonymous):

The equation to find the volume of a rectangle is LengthxWidthxHeight=Volume to solve back track. We already know what LengthxWidth is, as that is the formula to find the area of the rectangle which is 20cm^2 so now we just have to find the height. 20xheight=100, 100=the Volume (20xheight)/20=100/20 divide 20 from both sides height=5 so the answer is 5cm

OpenStudy (anonymous):

V = base * height ->h = V/ base = 100/ 20 = 5 cm